OpenStudy (anonymous):

1. Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation B: how many liters of orange juice are present in the can? show your work

OpenStudy (anonymous):

Part A: choose the variable and state what it represents: what percentage of water is in the can? create an equation for Part A: solve the equation: Part B: if the entire can holds 80% of orange juice, how many liter of orange juice are in the can?

OpenStudy (gorv):

let the total juice mixture= x litre juice=x*80/100=0.8x now in it water =mixture-juice=x-0.8x=0.2 x but water=1 litre x=1/0.2